OpenStudy (anonymous):

a weightles thread can bear upto 3.7kgwt a stone of mass 500 gm is tied to it and revolved in a circular path of radius 4m in a vertical plane. if g=10m/s^2, then the maximum angular velocity of stone can be???

OpenStudy (anonymous):

no the ans is 16m/sec

OpenStudy (anonymous):

Tension is maximum at the bottom point where T= mg+mrw^2 \[w = \sqrt{T - mg/mr}\] w = sqrt (37 -5 / 2)= 4 rad/s tangential velocity v =rw =16m/s

OpenStudy (anonymous):

the formula how it comes